OpenStudy (anonymous):

Solve the equation! √6x-1-2=0

OpenStudy (anonymous):

is that all under the radical?

OpenStudy (anonymous):

No the radical ends after the one.

OpenStudy (anonymous):

x=3/√6

OpenStudy (anonymous):

\[\sqrt{6x-1}-2=0\]

OpenStudy (anonymous):

oh. you could try moving the two to the other end of the equation, then squaring both sides

OpenStudy (anonymous):

Thats what it looks like.

OpenStudy (anonymous):

√6x-1-2=0 x=3/√6

OpenStudy (anonymous):

understand?

OpenStudy (anonymous):

That's the answer?

OpenStudy (anonymous):

x=3/√6

OpenStudy (anonymous):

Alright thanks!

OpenStudy (anonymous):

wait, what? I got 5/6